OpenStudy (anonymous):

derivative of y=ln sqrt(x^2+3)

OpenStudy (anonymous):

neophyte

OpenStudy (anonymous):

\[\ln \left(\sqrt{x^2+3}\right)\] same thing use chain rule, first take derivative of ln(), the sqrt[], and finally x^2+3

OpenStudy (anonymous):

\[\frac{x}{x^2+3} \]